OpenStudy (anonymous):

A car traveling on a highway is 60 miles away from home after 1 hour, and is 300 miles away from home after 5 hours. What is the rate of change of the car’s distance over time? A. mph B. mph C. 55 mph D. 60 mph

OpenStudy (anonymous):

Rate of change of car distance over time = (Change of distance) / (Change of Time) Change of distance = Final - Initial distance = 300 - 60 = 5 Change of Time = Final - Initial time = 5 - 1 = 4 Rate of change = 5 / 4